Circular No. 8647

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ma, 2005mb, 2005mc, 2005md, 2005me Additional supernovae have been discovered on unfiltered CCD im-

ages: 2005ma and 2005mb by T. Puckett (cf. IAUC 8646; with T. Orff for 2005ma and with T. Ceravolo for 2005mb); 2005mc by the Tsinghua Center for Astrophysics (THCA) Supernova Survey [F.-Y. Bian, Y.-Q. Lou, X.-F. Wang, and Y. Yang, THCA, Tsinghua University, Beijing; and Y.- L. Qiu, National Astronomical Observatories of China (NAOC)] with the 0.8-m THCA-NAOC Telescope (TNT); 2005md by W. Li (LOSS/KAIT; cf. IAUC 8646); and 2005me by L. A. G. Monard (cf. IAUC 8581).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005ma Dec. 24.21 4^{h}49^{m}53.^{s}91 −10^{o}45′23′′.4 17.9 1′′.9 W, 7′′.2 S
2005mb Dec. 21.46 13 05 52.46 +41 42 59.0 19.3 5′′.5 E, 20′′.2 S
2005mc Dec. 23.73 8 27 06.36 +21 38 45.6 17  4′′.8 E, 3′′.2 N
2005md Dec. 25.35 6 47 19.11 +33 34 24.9 18.1 21′′.8 E, 22′′.9 N
2005me Dec. 23.78 1 30 09.22 −42 41 07.0 17.5 41′′ E, 3′′ N

Additional magnitudes for 2005ma in MCG –02-13-13: Dec. 3, [19.3 (Puckett); 25.20 UT, 17.3 (J. Newton, Portal, AZ, 0.35-m reflector). Nothing was present at the location of 2005ma on numerous images taken by Puckett during 2000–2005 (limiting magnitudes ranging from 18.0 to 20.0). SN 2005ma has a spectrum similar to that of the type-IIn supernova 1997ab one year after discovery (cf. CBET 334). Additional magnitudes for 2005mb in NGC 4963: Dec. 5, [19.6 (Puckett); 22.41, 19.3 (Puckett); 24.71, R = 19.5 (H. Naito, Nishi-Harima Astronomical Observatory, 2.0-m Nayuta telescope). A measurement of Naito’s image by S. Kouzuma and H. Yamaoka, Kyushu University, yields position end figures 52^{s}.315, 59′′.43 (rms residuals ∼ 0′′.2) for 2005mb. Yamaoka adds that a Digitized Sky Survey (DSS) image taken on 1990 Jan. 25 shows nothing at the position of 2005mb (limiting red mag ∼ 20.5). Nothing was visible at the location of 2005mb on images taken by Puckett on numerous dates during 1998–2005 (limiting mag ranging from 18.7 to 20.2). Additional magnitudes for SN 2005mc in UGC 4414: Nov. 29, [18.5 (TNT); Dec. 5.43, [19.5 (Li, KAIT); 24 (in poor seeing) and 25.63, 16.5 (TNT). SN 2005mc is a type-Ia supernova (cf. CBETs 334, 338). Additional magnitudes for SN 2005md in NGC 2274: Dec. 10.40, [19.5 (KAIT); 25.86, 18.0 (X.-F. Wang and F.-Y. Bian, 0.8-m THCA-NAOC telescope, Xing-long Station). Additional magnitudes for SN 2005me in ESO 244-31: Dec. 2.93, [18.7; 23.806, 17.5; 26.760, 17.6. SN 2005me is not visible on the DSS (limiting red mag 20.5).

2005 December 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green